OpenStudy (anonymous):

34e + 36 = -9e + 337 E = ? @DJ♬Øϟϟ¥

OpenStudy (anonymous):

It's False :o

OpenStudy (anonymous):

What's false

OpenStudy (anonymous):

The answer.

OpenStudy (anonymous):

For what

OpenStudy (anonymous):

For this problem

OpenStudy (anonymous):

Why what did u get

OpenStudy (anonymous):

34e+36=-9e+337 subract 36 from both sides 34e=-9e+301 add -9e to both sides 25e=301 divide 25 both sides e=301/25 = 12.05

OpenStudy (anonymous):

Oh :o well, good job Shay! c:

OpenStudy (anonymous):

Glad to help! ^_^